# Conditioning of Banach Space Valued Gaussian Random Variables: An Approximation Approach Based on Martingales

### v1

@@ 3 Main Results
with convergence in $E$. Here, the last identity follows from part viii) of Theorem 3.4 applied to all $x^{\prime}\in\mathfrak{D}$ of a countable and $w^{*}$-dense subset $\mathfrak{D}$ of $B_{E^{\prime}}$, see Theorem 2.1, and a subsequent application of Hahn-Banach’s theorem. Moreover, part viii) of Theorem 3.4 also shows $C_{X|X}x^{\prime}=0$ for all $x^{\prime}\in E^{\prime}$, that is $C_{X|X}=0$, and hence we find
@@
for all $\omega\in\Omega\setminus N$. In other words, with increasing $n$, the distributions $\mathrm{P}_{X|X_{n}}(\,\cdot\,|X_{n}(\omega))$ concentrate more and more around the observation $X(\omega)$. Note that these convergence results can also be interpreted from a frequentist point of view for the Bayes method given by the posterior distributions $\mathrm{P}_{X|X_{n}}$ and the prior distribution $\mathrm{P}_{X}$. Here, it is assumed that the “true” parameter is some $X(\omega_{0})$. The convergence results in (18) and (19) then provide a contraction result in the sense of [16, Section 7.3], while (20) is an additional, yet interesting convergence.
4 Examples with a Focus on the Gaussian Process Case
Theorem 3.4 crucially depends on the existence and choice of filtering sequences. In this section, we elaborate on this by first showing that for separable Banach spaces $F$ we can always find a filtering sequence. In addition, we quickly present some natural filtering sequences for Hilbert spaces and reproducing kernel Hilbert spaces. Finally, we provide a detailed discussion for the Gaussian process case presented in the introduction, by considering $E:=C(T)$ and $F:=C(S)$ for a closed $S\subset T$.
Our first result shows the existence of a filtering sequence for generic, separable $F$, where we recall from Theorem 2.1 that for such $F$ we always have a countable and $w^{*}$-dense $\mathfrak{D}_{F^{\prime}}\subset B_{F^{\prime}}$,
@@ 4 Examples with a Focus on the Gaussian Process Case
Combining Proposition 4.1 with Theorem 3.4 we see, for example, that the regular conditional probabilities $\mathrm{P}_{X|Y}$ of jointly Gaussian random variables are always almost surely Gaussian measures. Moreover, we can always approximate their means and covariances by the described finite dimensional projections.
We have already seen in Proposition 3.3 that sequences $(y_{i}^{\prime})\subset F^{\prime}$ whose span is $w^{*}$-sequentially dense in $F^{\prime}$ can be used to construct filtering sequences. The following result provides a characterization of such sequences that can be easier verified in some situations.
@@
Let us now turn to more specific sequences. The following lemma considers orthonormal bases (ONBs) in separable Hilbert spaces.
Lemma 4.3.
@@
If $H$ is a separable reproducing kernel Hilbert space (RKHS) on a set $T$ with kernel $k:T\times T\to\mathbb{R}$, see e.g. [38, Chapter 4.2], then we can, of course, apply Lemma 4.3 to construct a filtering sequence. However, for RKHSs we have an additional generic way to construct a filtering sequence. To present this, let us recall, that $\operatorname{d}_{k}:T\times T\to[0,\infty)$ defined by
@@
gives a pseudo-metric on $T$ called the kernel metric, see e.g. [38, Chapter 4.3]. With this preparation, the RKHS case reads as follows.
@@
Let $H$ be a separable RKHS on $T$ with kernel $k$. Then there exists a countable $T_{0}\subset T$ that is dense in $T$ with respect to $\operatorname{d}_{k}$. Moreover, if we have a countable $T_{0}=\{t_{i}^{\prime}:i\geq 1\}$ that is $\operatorname{d}_{k}$-dense in $T$, then $A_{n}:H\to\mathbb{R}^{n}$ defined by
$A_{n}(y):=\bigl{(}\bigl{\langle}k(t_{1},\,\cdot\,),y\bigr{\rangle}_{H},\dots,\bigl{\langle}k(t_{n},\,\cdot\,),y\bigr{\rangle}_{H}\bigr{)}\,,\qquad\qquad y\in H$
gives filtering sequence $(A_{n})$ for $H$.
In most cases, working with the kernel metric may be somewhat cumbersome. For this reason we note that if $\operatorname{d}$ is a metric on $T$ such that $k$ is continuous with respect to $\operatorname{d}$, then $\operatorname{d}$-convergence implies $\operatorname{d}_{k}$-convergence, see e.g. [38, Lemma 4.29]. Consequently, if $(T,\operatorname{d})$ is separable, and $T_{0}$ is a countable and $\operatorname{d}$-dense subset of $T$, then it is also a $\operatorname{d}_{k}$-dense subset of $T$. In other words, we may replace $\operatorname{d}_{k}$ by $\operatorname{d}$ in Proposition 4.4.
Our next goal is to consider Gaussian random variables on the Banach space $C(T)$ of continuous functions, where $(T,\operatorname{d})$ is a compact metric space. To this end, we first note that in this case, $C(T)$ is a separable, see e.g. [11, Theorem V.6.6] or [14, Corollary 11.2.5]. To describe the dual of $C(T)$, we write ${\cal M}(T)$ for the space of all finite, signed Borel measures equipped with the total variation norm. Riesz’ representation theorem then shows that
@@
is an isometric isomorphism, see e.g. [10, Theorem 7.3.6] and [4, Corollary 29.13], where for the first reference we note that compact metric spaces are Polish and Borel measures on Polish spaces are regular, and for the second reference we note that compact metric spaces have a countable base. In particular, the Dirac measures are mapped to the point the evaluation functionals $\delta_{t}:C(T)\to\mathbb{R}$ defined by $\delta_{t}f=f(t)$ for $t\in T$ and $f\in C(T)$.
@@
It is well-known that a Gaussian process $(X_{t})_{t\in T}$ with continuous paths can be described by the mean and covariance functions given by (1) and (2). If $X$ is the Gaussian random variable associated to the process via (23), then the mean $\mathbb{E}X$ and the covariance operator $\operatorname{cov}(X)$ can be used to compute the mean and covariance function by
@@
$k(s,t)$ $=\langle\delta_{s},\operatorname{cov}(X)\delta_{t}\rangle_{C(T)^{\prime},C(T)}$ (25)
for all $s,t\in T$, where we have used (A.11), respectively (B.1). Conversely, if we have the mean function $m$, then the identity above shows that we known the function $\mathbb{E}X\in C(T)$ pointwise, i.e. we know $\mathbb{E}X$. Moreover, the covariance function determines the covariance operator, namely, one can show
@@
for the distribution of a given Gaussian random variable $X:\Omega\to C(T)$, where $m$ and $k$ are mean and covariance function of the associated stochastic process.
In summary, instead of working with a Gaussian random variable $X:\Omega\to C(T)$ we can alternatively consider Gaussian processes $(X_{t})_{t\in T}$ with continuous paths together with their mean and covariance function. In the following we will often adopt this point of view, which is common in the Gaussian process for machine learning literature, see e.g. [36, 24]. In particular, we speak of conditioned processes, when we condition the associated Gaussian random variables.
Let us now assume that we have a Gaussian processes $(X_{t})_{t\in T}$ with continuous paths and that we have an observation $(Y_{s})_{s\in S}$ of this process on a non-empty subset $S\subset T$. In other words, we have $Y_{s}(\omega)=X_{s}(\omega)$ for all $s\in S$ and $\omega\in\Omega$, where $(\Omega,{\cal A},\mathrm{P})$ is the underlying probability space and $(T,\operatorname{d})$ is a compact metric space. For technical simplicity, we assume that $S$ is closed, and thus also a compact metric space. To describe this setup by Gaussian random variables, we write $E:=C(T)$ and $F:=C(S)$ and denote the Gaussian random variables associated to the two processes above via (23) by $X:\Omega\to C(T)$ and $Y:\Omega\to C(S)$. This gives
@@
that is $Y=\,\cdot\,_{|S}\circ X$, where the restriction operator $\,\cdot\,_{|S}:C(T)\to C(S)$ is bounded and continuous.
We first consider the case of a finite set $S=\{s_{1},\dots,s_{n}\}$. Here we write $I_{S}:C(T)\to\mathbb{R}^{n}$ for the bounded linear operator defined by
@@
Clearly, the general setup (26) reduces to $Y=I_{S}\circ X$ as $C(S)$ can be identified by $\mathbb{R}^{n}$. Finally, note that if we have another finite set $R=\{r_{1},\dots,r_{m}\}$, then the covariance operator $\operatorname{cov}(I_{R}\circ X,I_{S}\circ X):\mathbb{R}^{n}\to\mathbb{R}^{m}$ can be represented by the $m\times n$-matrix
@@
With these preparations, the application of Theorem 3.1 to the observational model (26) reads as follows.
@@
Before we proceed with conditioning on observations on infinite $S$, we note that we can easily modify the observational model (28) to e.g.
@@
for all $t,t_{1},t_{2}\in T$. Not surprisingly, these formulas are the well-known textbook formulas from the Gaussian processes for machine learning literature, see e.g. [36]. However, there is is subtle but important difference in their meaning: The mentioned textbook formulas do not ensure that a Gaussian process with the above mean and covariance function has a version with continuous paths and even if this can be derived by other means, they do not ensure that the resulting measure of such a version on $C(T)$ equals the regular conditional measure we seek. In contrast, our approach establishes exactly these guarantees, so that the common functional interpretation of Gaussian processes for machine learning are indeed justified. This is similar in spirit of [45, Corollary 5], where more general functionals on $C(T)$ are considered in the noise free case $\varepsilon=0$.
Let us now turn to observations on infinite $S$. To apply Theorem 3.4 for this setting, we need a filtering sequence since the $C(S)$-valued random variable $Y:=X_{|S}$ lives in an infinite dimensional space. The following proposition provides a natural filtering sequence.
Proposition 4.7.
Let $(T,\operatorname{d})$ be a compact metric space and $T_{\infty}\subset T$ be a countable, dense subset with enumeration $T_{\infty}=\{t_{i}:i\geq 1\}$. For $n\geq 1$ we consider the evaluation operator $I_{n}:C(T)\to\mathbb{R}^{n}$ defined by
$I_{n}(f):=\bigl{(}f(t_{1}),\dots,f(t_{n})\bigr{)}\,,\qquad\qquad f\in C(T)$
Then $(I_{n})$ is a filtering sequence for $C(T)$.
Let us consider a closed subset $S\subset T$ of a compact metric space and a countable, dense $S_{\infty}\subset S$ with enumeration $S_{\infty}=\{s_{i}:i\geq 1\}$. Then one can view the filtering sequence $(I_{n})$ of Proposition 4.7 as a mean to approximate the infinite dimensional observational random variable $Y:=X_{|S}$ with the help of the sequence $(Y_{n})$ of finite dimensional observations $Y_{n}:=X_{|S_{n}}$ taken at the points $S_{n}:=\{s_{1},\dots,s_{n}\}$. In other words, we approximate $Y$ on an arbitrarily fine, finite set of observational points. Since we have $Y_{n}=I_{n}\circ X=I_{S_{n}}\circ X$ the combination of Theorem 3.4 with Theorem 4.6 is straightforward and reads as follows.
@@
Let $(\Omega,{\cal A},\mathrm{P})$ be a probability space, $(T,\operatorname{d})$ be a compact metric space, and $(X_{t})_{t\in T}$ be a Gaussian processes on $(\Omega,{\cal A},\mathrm{P})$ with continuous paths. Moreover, let $X:\Omega\to C(T)$ be the Gaussian random variable associated to the process via (23), $S\subset T$ be closed, and
@@
Finally, let $S_{\infty}\subset S$ be countable and dense with enumeration $S_{\infty}=\{s_{i}:i\geq 1\}$. We write $S_{n}:=\{s_{1},\dots,s_{n}\}$ for all $n\geq 1$. Then for every version $\mathrm{P}_{X|Y}(\,\cdot\,|\,\cdot\,):\mathfrak{B}(C(T))\times C(S)\to[0,1]$ of regular conditional probability of $X$ given $Y$ there exist an $N\in\mathfrak{B}(C(S))$ with $\mathrm{P}_{Y}(N)=0$ such that for all $g\in C(S)\setminus N$ we have
@@
where the convergence is uniform in $t\in T$, respectively in $t_{1},t_{2}\in T$. Moreover, for all $g\in C(S)\setminus N$, $s\in S$, and $t\in T$ we have
@@
Of course, the remaining assertions of Theorem 3.4 also hold in the special setup considered in Theorem 4.8. In particular note that the convergence of the covariance operators in Theorem 3.4 are with respect to the nuclear norm, while the uniform convergence of the covariance functions stated in Theorem 4.8 is a consequence of the weaker operator norm convergence. In addition, it is, of course, possible to translate Theorems 3.5 and 3.6 to the $C(T)$-valued case with the help of Lemma 6.2.

### v3

@@ 3 Main Results
with convergence in $E$. Here, in the last identity we used viii) of Theorem 3.3 together with the simple observation that $\langle x^{\prime},X\rangle_{E^{\prime},E}:\Omega\to\mathbb{R}$ is $\sigma(X)$-measurable for all $x^{\prime}\in E^{\prime}$, and subsequently we applied Hahn-Banach’s theorem. Similarly, part viii) of Theorem 3.3 also shows $C_{X|X}x^{\prime}=0$ for all $x^{\prime}\in E^{\prime}$, that is $C_{X|X}=0$, and hence we find
@@
for all $\omega\in\Omega\setminus N$. In other words, with increasing $n$, the distributions $\mathrm{P}_{X|X_{n}}(\,\cdot\,|X_{n}(\omega))$ concentrate more and more around the observation $X(\omega)$. Note that these convergence results can also be interpreted from a frequentist point of view for the Bayes method given by the posterior distributions $\mathrm{P}_{X|X_{n}}$ and the prior distribution $\mathrm{P}_{X}$, if it is assumed that the “true” parameter is some $X(\omega_{0})$: Namely, the convergence results in (16) and (17) then provide a contraction result in the sense of [19, Section 7.3], while (18) is an additional convergence.
4 Filtering Sequences
Theorems 3.3 and 3.5 crucially depend on the existence of filtering sequences and our ability to construct and apply them. In this section, we therefore investigate filtering sequences in some detail. In particular, we show that for separable Banach spaces $F$ there always exists a filtering sequence. In addition, we present some natural and readily available filtering sequences for Hilbert spaces and reproducing kernel Hilbert spaces, as well as the space of continuous functions over some compact metric space. Finally, we also investigate general Banach spaces consisting of functions.
Our first result shows that a sufficiently rich sequence $(y_{i}^{\prime})\subset F^{\prime}$ can be used to construct a filtering sequence. We will see below that the key advantage of this proposition is that it translates the measure-theoretic definition of filtering sequences into the language of functional analysis.
@@ 4 Filtering Sequences
Then $(A_{n})$ is a filtering sequence for $F$.
Note that (19) is, for example, satisfied, if we have a sequence $(y_{i}^{\prime})\subset F^{\prime}$, for which (19) even holds for the norm-closure, that is,
$\overline{\operatorname{span}\{y_{i}^{\prime}:i\geq 1\}}^{\|\cdot\|_{F^{\prime}}}=F^{\prime}\,.$ (21)
Unfortunately, a sequence $(y_{i}^{\prime})\subset F^{\prime}$ with (21) exists if and only if $F^{\prime}$ is separable. As a consequence, spaces such as $F:=C([0,1]^{d})$ do not enjoy (21), and for this reason we will mostly focus on (19).
Now, our first main result of this section shows the existence of a filtering sequence for generic, separable $F$, where we recall from Theorem 2.1 that for such $F$ we always have a countable and $w^{*}$-dense $\mathfrak{D}_{F^{\prime}}\subset B_{F^{\prime}}$.
Theorem 4.2.
@@
Combining Theorem 4.2 with Theorem 3.3 we see, for example, that the regular conditional probabilities $\mathrm{P}_{X|Y}$ of jointly Gaussian random variables $X$ and $Y$ are always almost surely Gaussian measures. Moreover, we can always approximate their means and covariances by the described finite dimensional projections.
We have already seen in Proposition 4.1 that sequences $(y_{i}^{\prime})\subset F^{\prime}$ whose span is $w^{*}$-sequentially dense in $F^{\prime}$ can be used to construct filtering sequences. Unfortunately, working with the $w^{*}$-sequential closure in (19) is often rather complicated. For this reason the following result provides a characterization of sequences satisfying (19) that can be easier verified in some situations.
@@
Let us now turn to more specific spaces. The following corollary considers orthonormal bases (ONBs) in separable Hilbert spaces.
Corollary 4.4.
@@
Of course, it is possible to generalize Corollary 4.4 to Banach spaces $F$ whose dual $F^{\prime}$ have a Schauder base. We skip the rather obvious details.
We have seen in the introduction that one of our goals is to better understand GP4MLs. Since GP4MLs generate random functions we now investigate filtering sequences for Banach spaces consisting of functions (BSF). To this end, we say that a Banach space $F$ is a BSF on some set $T$, if $F$ consists of functions $T\to\mathbb{R}$, and the point evaluations $\delta_{t}:F\to\mathbb{R}$ given by $\delta_{t}f=f(t)$ are continuous for all $t\in T$.
Note that BSFs span a wide class of spaces. For example, the space $\ell_{\infty}(T)$ of all bounded functions $T\to\mathbb{R}$ equipped with the supremums norm $\|\cdot\|_{\infty}$ is a BSF. In addition, the space $C_{b}(T,\operatorname{d})$ of bounded, continuous functions $T\to\mathbb{R}$ on some (pseudo) metric space $(T,\operatorname{d})$ is a BSF, if we again equip $C_{b}(T,\operatorname{d})$ with the supremums norm. Moreover, reproducing kernel Hilbert spaces (RKHSs) are covered, as they are by definition both Hilbert spaces and BSFs, see e.g. [42, Chapter 4.2] for details. Furthermore, spaces of bounded, $\alpha$-Hölder continuous functions as well as spaces of bounded, continuously differentiable functions are BSFs. In addition, Sobolev spaces can be viewed as BSFs if their smoothness is sufficiently large in the sense that Sobolev’s embedding theorem is satisfied. Since BSFs have recently attracted some interest in the machine learning community, we refer for more information on BSFs to e.g. [4, 41] and the references mentioned therein. Finally, note that BSFs also play an important role for the PDE applications mentioned in the introduction, see e.g. [38].
Now, if we have BSF $F$ on $T$ then it is somewhat natural to use point evaluations to construct filtering sequences. Namely, if we have some $S:=\{t_{1},\dots,t_{n}\}\subset T$, then $I_{S}:F\to\mathbb{R}^{n}$ defined by
$I_{S}(f):=\bigl{(}\langle\delta_{t_{1}},f\rangle_{F^{\prime},F},\dots,\langle\delta_{t_{n}},f\rangle_{F^{\prime},F}\bigr{)}\,,\qquad\qquad f\in F$ (22)
is a bounded linear operator. The following result shows that in certain situations these operators can be used to construct filtering sequences.
Theorem 4.5.
Let $(T,\operatorname{d})$ be a separable pseudo-metric space and $F$ be a separable BSF on $T$ for which all $f\in F$ are continuous. Furthermore, let $T_{\infty}\subset T$ be a countable and dense subset with enumeration $T_{\infty}=\{t_{i}:i\geq 1\}$. Then $(\delta_{t_{i}})\subset F^{\prime}$ is separating and $(I_{T_{n}})$ is a filtering sequence for $F$, where we have set $T_{n}:=\{t_{1},\dots,t_{n}\}$.
Recall that compact metric spaces $(T,\operatorname{d})$ are separable and that in this case the corresponding space $C(T)$ of continuous functions $f:T\to\mathbb{R}$ is also separable, e.g. [13, Theorem V.6.6] or [16, Corollary 11.2.5]. Therefore, we obtain the following corollary of Theorem 4.5 without any further proof.
Corollary 4.6.
Let $(T,\operatorname{d})$ be a compact metric space and $T_{\infty}\subset T$ be a countable, dense subset with enumeration $T_{\infty}=\{t_{i}:i\geq 1\}$. Then $(I_{T_{n}})$ is a filtering sequence for $C(T)$, where we have set $T_{n}:=\{t_{1},\dots,t_{n}\}$.
In the next result we will consider the space $C^{1}([0,1])$ of continuous functions $f:[0,1]\to\mathbb{R}$ that are continuously differentiable on $(0,1)$ and for which there exists a continuous $g:[0,1]\to\mathbb{R}$ with $g_{|(0,1)}=f^{\prime}$. Moreover, we equip this space with the norm
$\|f\|_{C^{1}([0,1])}:=\|f\|_{\infty}+\|f^{\prime}\|_{\infty}\,,\qquad\qquad f\in C^{1}([0,1]).$
It is well-known that $C^{1}([0,1])$ is a separable Banach space and from this we quickly conclude that it is also a BSF. Now, the following simple lemma provides a separating sequence for $C^{1}([0,1])$ and thus also a filtering sequence via Proposition 4.3 and Proposition 4.1.
Lemma 4.7.
Let $\{t_{i}:i\geq 1\}\subset(0,1)$ be dense in $[0,1]$ and $y_{i}^{\prime}:C^{1}([0,1])\to\mathbb{R}$ be defined by $y_{i}^{\prime}(f):=f^{\prime}(t_{i})$ for all $f\in C^{1}([0,1])$. If we further set $y_{0}^{\prime}:=\delta_{0}$, then $(y_{i}^{\prime})_{i\geq 0}$ is a separating sequence for $C^{1}([0,1])$.
Note that in view of ODE applications mentioned in the introduction, Lemma 4.7 provides an interesting alternative to the point evaluations considered in Corollary 4.6. Moreover, generalizations of Lemma 4.7 to higher dimensions and derivatives are straightforward.
Our next goal is to apply Theorem 4.5 to RKHSs. To this end, recall that every RKHS $H$ on $T$ has a unique kernel $k:T\times T\to\mathbb{R}$, which by the reproducing property can be used to represent the point evaluations. More precisely, we have
$\langle\delta_{t},h\rangle_{H^{\prime},H}=\langle k(t,\,\cdot\,),h\rangle_{H}\,,\qquad\qquad t\in T,\,h\in H,$
e.g. [42, Chapter 4.2]. With these preparations, the application of Theorem 4.5 to RKHSs reads as follows.
Corollary 4.8.
Let $(T,\operatorname{d})$ be a separable, pseudo-metric space and $H$ be an RKHS on $T$ with continuous kernel $k:T\times T\to\mathbb{R}$. Then $H$ is separable and for all countable and dense subsets $T_{\infty}=\{t_{i}:i\geq 1\}$ of $T$ the operators $I_{n}:H\to\mathbb{R}^{n}$ defined by
$I_{n}(h):=\bigl{(}\langle k(t_{1},\,\cdot\,),h\rangle_{H},\dots,\langle k(t_{n},\,\cdot\,),h\rangle_{H}\bigr{)}\,,\qquad\qquad h\in H$ (23)
give a filtering sequence $(I_{n})$ for $H$.
In the rest of this section, we show that for all separable BSFs $F$ on $T$ it is possible to construct a pseudo metric on $T$ such that the assumptions of Theorem 4.5 are satisfied.
We first consider the case of separable RKHSs $H$ on some set $T$ since in this case, the construction is straightforward. To begin with, let us recall, that $\operatorname{d}_{k}:T\times T\to[0,\infty)$ defined by
@@
gives a pseudo-metric on $T$ called the kernel metric, see e.g. [42, Chapter 4.3]. With this preparation, the RKHS case reads as follows.
@@
Let $H$ be a separable RKHS on $T$ with kernel $k$. Then $(T,\operatorname{d}_{k})$ is a separable pseudo-metric space and all $h\in H$ are continuous with respect to $\operatorname{d}_{k}$.
Combining Proposition 4.9 with Theorem 4.5 we see that (23) defines a filtering sequence if we choose a countable and $\operatorname{d}_{k}$-dense subset of $(T,\operatorname{d}_{k})$.
Our last result generalizes Proposition 4.9 to separable BSFs $F$. Here, a natural idea is to pull back a metric from $F^{\prime}$ to $T$ with the help of the “feature map” $\Phi:T\to F^{\prime}$ given by $t\mapsto\delta_{t}$ as we have done for the kernel metric. Unfortunately, however, we can no longer take the norm of $F^{\prime}$ for this approach as the needed separability of $T$ would require $F^{\prime}$ to be separable. Moreover, simply taking a metric describing the $w^{*}$-topology on a (scaled) ball of $B_{F^{\prime}}$ is not possible either, as the above $\Phi$ is bounded if and only if $F\subset\ell_{\infty}(T)$, and the latter is an additional assumption we would like to avoid. Nonetheless, the general idea is possible if we consider a slightly modified map and a different metric, as the proof of the following theorem reveals.
Theorem 4.10.
Let $F$ be a separable BSF on $T$. Then there exists a pseudo-metric $\operatorname{d}$ on $T$ such that $(T,\operatorname{d})$ is separable and all $f\in F$ are continuous with respect to $\operatorname{d}$. In addition, there exists a countable subset $T_{\infty}=\{t_{i}:i\geq 1\}$ of $T$ such that the sequence $(\delta_{t_{i}})\subset F^{\prime}$ is separating.
Let us assume for a moment that $F$ is a separable BSF on $T$ and $T_{\infty}=\{t_{i}:i\geq 1\}$ is a set as in Theorem 4.10. Then Proposition 4.3 gives
$\overline{\operatorname{span}\{\delta_{t_{i}}^{\prime}:i\geq 1\}}^{w^{*}\mathrm{\!-seq}}=F^{\prime}\,,$
and Theorem 4.5 ensures that $(I_{T_{n}})$ is a filtering sequence for $F$, where we have set $T_{n}:=\{t_{1},\dots,t_{n}\}$. In other words, for separable BSFs it is always possible to apply our main Theorems 3.3 and 3.5 with filtering sequences based upon point evaluations.
5 Conditioning Gaussian Processes with Continuous Paths
In this section we illustrate how the general results of Section 3 can be applied to the GP4ML setting. Namely, we will consider the conditioning of Gaussian random variables for the spaces $E:=C(T)$ and $F:=C(S)$, where $(T,\operatorname{d})$ is a compact metric space and $S\subset T$ is closed. In addition, we will discuss how these results can be formulated from a Gaussian process perspective.
To apply our general results we first note that $C(T)$ is a separable, see e.g. [13, Theorem V.6.6] or [16, Corollary 11.2.5]. Moreover, to describe the dual of $C(T)$, we write ${\cal M}(T)$ for the space of all finite, signed Borel measures equipped with the total variation norm. Riesz’ representation theorem then shows that
@@ 5 Conditioning Gaussian Processes with Continuous Paths
is an isometric isomorphism, see e.g. [12, Theorem 7.3.6] and [6, Corollary 29.13], where for the first reference we note that compact metric spaces are Polish and Borel measures on Polish spaces are regular, and for the second reference we note that compact metric spaces have a countable base. Finally note that (24) maps the Dirac measures $\delta_{\{t\}}$ to the point evaluation functionals $\delta_{t}:C(T)\to\mathbb{R}$, that is $\hat{\delta}_{\{t\}}=\delta_{t}$ for all $t\in T$.
@@
It is well-known that a Gaussian process $(X_{t})_{t\in T}$ with continuous paths can be described by the mean and covariance functions given by (1) and (2). Moreover, if $X$ is the Gaussian random variable associated to the process via (26), then the mean $\mathbb{E}X$ and the covariance operator $\operatorname{cov}(X)$ can be used to compute the mean and covariance function, namely we have
@@
$k(t_{1},t_{2})$ $=\langle\delta_{t_{1}},\operatorname{cov}(X)\delta_{t_{2}}\rangle_{C(T)^{\prime},C(T)}$ (28)
for all $t,t_{1},t_{2}\in T$, where we have used (A.11), respectively (B.1). Conversely, if we have the mean function $m$, then the identity above shows that we known the function $\mathbb{E}X\in C(T)$ pointwise, i.e. we know $\mathbb{E}X$. Moreover, the covariance function determines the covariance operator, namely, one can show
@@
for the distribution of a given Gaussian random variable $X:\Omega\to C(T)$, where $m$ and $k$ are mean and covariance function of the associated Gaussian process.
In summary, instead of working with a Gaussian random variable $X:\Omega\to C(T)$ we can alternatively consider Gaussian processes $(X_{t})_{t\in T}$ with continuous paths together with their mean and covariance function. In the following we will often adopt this point of view, which is common in the Gaussian process for machine learning literature, see e.g. [7, 39, 27]. In particular, we speak of conditioned processes, when we condition the associated Gaussian random variables.
Let us now assume that we have a compact metric space $(T,\operatorname{d})$, a Gaussian processes $(X_{t})_{t\in T}$ with continuous paths, and an observation $(Y_{s})_{s\in S}$ of this process on a non-empty subset $S\subset T$. In other words, we have $Y_{s}(\omega)=X_{s}(\omega)$ for all $s\in S$ and $\omega\in\Omega$, where $(\Omega,{\cal A},\mathrm{P})$ is the underlying probability space. For technical simplicity, we assume that $S$ is closed, and thus also a compact metric space. To describe this setup by Gaussian random variables, we write $E:=C(T)$ and $F:=C(S)$ and denote the Gaussian random variables associated to the two processes above via (26) by $X:\Omega\to C(T)$ and $Y:\Omega\to C(S)$. This gives
@@
that is $Y=\,\cdot\,_{|S}\circ X$, where the restriction operator $\,\cdot\,_{|S}:C(T)\to C(S)$ is bounded and linear.
We first consider the case of a finite set $S=\{s_{1},\dots,s_{n}\}$. Here we recall from (22) that the bounded linear operator $I_{S}:C(T)\to\mathbb{R}^{n}$ is defined by
@@
Clearly, the general setup (29) reduces to $Y=I_{S}\circ X$ as $C(S)$ can be canonically identified by $\mathbb{R}^{n}$. Finally, note that if we have another finite set $R=\{r_{1},\dots,r_{m}\}$, then the covariance operator $\operatorname{cov}(I_{R}\circ X,I_{S}\circ X):\mathbb{R}^{n}\to\mathbb{R}^{m}$ can be represented by the $m\times n$-matrix
@@
With these preparations, the application of Theorem 3.1 to the specific observational model (29) reads as follows.
@@
Before we proceed with conditioning on observations on infinite $S$, we note that we can easily generalize the observational model (31) to e.g.
@@
for all $t,t_{1},t_{2}\in T$. Not surprisingly, these formulas are the well-known textbook formulas from the GP4MLs literature, see e.g. [39, Chapter 2.2]. However, there is is subtle but important difference in their meaning: The mentioned textbook formulas do not ensure that a Gaussian process with the above mean and covariance function has a version with continuous paths and even if this can be derived by other means, they do not ensure that the resulting measure of such a version on $C(T)$ equals the regular conditional probability we seek. In contrast, our approach establishes exactly these guarantees, so that the common functional interpretation of GP4MLs is indeed justified. This is similar in spirit to [49, Corollary 5], where more general functionals on $C(T)$ are considered in the noise free case $\varepsilon=0$.
Let us now turn to observations on infinite, closed $S\subset T$, where again $(T,\operatorname{d})$ is a compact metric space. Since in this case $(S,\operatorname{d})$ is also a compact metric space, there exists a countable, dense $S_{\infty}\subset S$ with enumeration $S_{\infty}=\{s_{i}:i\geq 1\}$. Moreover, we have seen in Corollary 4.6 that the evaluation operators $I_{n}:C(S)\to\mathbb{R}^{n}$ defined by
$I_{n}(g):=\bigl{(}g(s_{1}),\dots,g(s_{n})\bigr{)}\,,\qquad\qquad g\in C(S)$
give a filtering sequence $(I_{n})$ for $C(S)$. Roughly speaking, this sequence provides us with a mean to approximate the infinite dimensional observational random variable $Y:=X_{|S}$ with the help of the sequence $(Y_{n})$ of finite dimensional observations $Y_{n}:=X_{|S_{n}}$ taken at the points $S_{n}:=\{s_{1},\dots,s_{n}\}$. In other words, we approximate $Y$ on an arbitrarily fine but finite set of observational points. Since we have $Y_{n}=I_{n}\circ X=I_{S_{n}}\circ X$ the combination of Theorem 3.3 with Theorem 5.2 is straightforward and reads as follows.
@@
Let $(\Omega,{\cal A},\mathrm{P})$ be a probability space, $(T,\operatorname{d})$ be a compact metric space, and $(X_{t})_{t\in T}$ be a Gaussian process on $(\Omega,{\cal A},\mathrm{P})$ with continuous paths. Moreover, let $X:\Omega\to C(T)$ be the Gaussian random variable associated to the process via (26), $S\subset T$ be closed, and
@@
Finally, let $S_{\infty}\subset S$ be countable and dense with enumeration $S_{\infty}=\{s_{i}:i\geq 1\}$. We write $S_{n}:=\{s_{1},\dots,s_{n}\}$ for all $n\geq 1$. Then for every version $\mathrm{P}_{X|Y}(\,\cdot\,|\,\cdot\,):\mathfrak{B}(C(T))\times C(S)\to[0,1]$ of the regular conditional probability of $X$ given $Y$ there exist an $N\in\mathfrak{B}(C(S))$ with $\mathrm{P}_{Y}(N)=0$ such that for all $g\in C(S)\setminus N$ we have
@@
and the convergence is uniform in $t\in T$, respectively in $t_{1},t_{2}\in T$. Moreover, for all $g\in C(S)\setminus N$, $s\in S$, and $t\in T$ we have
@@
Of course, the remaining assertions of Theorem 3.3 also hold in the special setup considered in Theorem 5.3. For example, the weak convergence $\mathrm{P}_{X|(X_{|S_{n}})}(\,\cdot\,|g(S_{n}))\to\mathrm{P}_{X|(X_{|S})}(\,\cdot\,|g)$ holds true. Also note that the convergence of the covariance operators in Theorem 3.3 are with respect to the nuclear norm, while the uniform convergence of the covariance functions stated in Theorem 5.3 is a consequence of the weaker operator norm convergence. Finally, it is, of course, possible to translate Theorems 3.4 and 3.5 to the $C(T)$-valued case, too.